OpenStudy (anonymous):

Sammy says that multiplying 4 1/8 by 3 and then subtracting 1 gives an answer of 11 3/8 explain how rounding can be used to decide if Sammy answer is right. I dont understand this PLEASE HELP ME!!!

OpenStudy (anonymous):

Here we go....\[4\frac{1}{8}=\frac{33}{8}\implies 3(\frac{33}{8})=\frac{99}{8}~so~\frac{99}{8}-1=\frac{91}{8}=11\frac{3}{8}\]